# Membrane Paradigm Approach to the Johannsen-Psaltis Black Hole

### v1

@@ III.3 The latest version of Johannsen-Psaltis Spacetime
The black hole, introduced in [9] suffers from a violation of the strong rigidity theorem, and carries a chaotic geodesic equation [23]. Its event horizon equation is given by a quintic equation which is not solvable by radicals. In [9], instead of the event horizon, the authors studied the Killing horizon which is not the same thing when strong rigidity is violated. These problems make it hard for us to study its transport coefficients in the limit of the true horizon. In any case, we give the membrane paradigm of that metric in the appendix. In [24], the authors reconsidered the metric given in [10] and made some judicious choices of the metric functions. The new metric in [24] violates the strong rigidity however, it obeys the roto-surface theorem and the weak rigidity theorem which allow one to analyze the black hole angular momentum and surface gravity while equating the event horizon to the Killing horizon. In the latest work [25], the authors considered the form suggested in [10]. This will be the metric we shall study here given in local coordinates as:
@@
While choosing the metric functions as:
@@
the metric deformations are kept dimensionless and for weak field tests, one needs to set $\epsilon_{2}=\alpha_{2}=0$. This form of the metric is asymptotically flat, reproduces the Newtonian effects in the limit and satisfies the PPN constraints. We will consider only the lowest-order deformation of the metric, not the whole summation of the deformation parameters. One can clearly, understand that the main difference between this metric and the metric in [9] is that it has more hair introduced via deformations however, it still has the Kerr event radius that is $g^{rr}|_{r_{H}}=0$ with $r_{H}=m+\sqrt{m^{2}+a^{2}}$. In our analysis, we should consider 3 different aspects:
@@
    We will fix the bulk viscosity of the rotating solution to be the same as the non-rotating one, i.e. $\zeta=-\frac{1}{16\pi}$.
@@
Using the metric functions, the generic metric can be recast as:
@@
Now, we should decompose this metric in the $(2+1+1)$-form:
@@
Let $u_{\mu}dx^{\mu}=F_{t}dt+\omega d\phi$, $n_{\mu}dx^{\mu}=F_{r}dr$. The structure of this metric can be put in the form of (44) such that one can directly start to calculate the factors relevant to the membrane description. The components of the extrinsic curvature are found to be :
@@
$K_{\theta\theta}$ $=\frac{\partial_{r}S}{2\sqrt{\frac{S}{\Delta Z}}},\quad K_{\phi\phi}=\frac{1}{E(r,\theta)}\left(\sin^{2}\theta S\left(\left(a^{2}+r^{2}\right)^{2}A^{2}-a^{2}\sin^{2}\theta\Delta\right)\right),\quad K_{rr}=0,$
where we introduced the functions:
@@
Following the construction of a generic Kerr-like Membrane Paradigm algorithm, one can find the other transport coefficients. For instance, the non-zero components of the shear tensor turn out to be
@@
and the non-zero component of the momentum becomes:
@@
Now, let us analyze what happens to the pressure in different limits of the deformation parameters. When the metric coefficients are plugged to (55) it becomes:

### v2

@@ III.3 The latest version of Johannsen-Psaltis spacetime
The black hole introduced in [9] suffers from a violation of the strong rigidity theorem, and carries a chaotic geodesic equation [23]. Its event horizon equation is given by a quintic equation that is not solvable by radicals. In [9], instead of the event horizon, the authors studied the Killing horizon which is not the same thing when strong rigidity is violated. These problems make it hard for us to study its transport coefficients in the limit of the true horizon. In [24], the authors reconsidered the metric given in [10] and made some judicious choices of the metric functions. The new metric in [24] violates the strong rigidity; however, it obeys the rotosurface theorem and the weak rigidity theorem that allow one to analyze the black hole angular momentum and surface gravity while equating the event horizon to the Killing horizon. In the latest work [25], the authors considered the form suggested in [10]. This will be the metric we shall study here given in local coordinates as
@@
While choosing the metric functions as
@@
the metric deformations are kept dimensionless and for weak field tests, one needs to set $\epsilon_{2}=\alpha_{2}=0$. This form of the metric is asymptotically flat, reproduces the Newtonian effects in the limit and satisfies the PPN constraints. We will consider only the lowest-order deformation of the metric, not the whole summation of the deformation parameters. One can clearly understand that the main difference between this metric and the metric in [9] is that it has more hair introduced via deformations; however, it still has the Kerr event radius, that is, $g^{rr}|_{r_{H}}=0$ with $r_{H}=m+\sqrt{m^{2}+a^{2}}$. In our analysis, we should consider three different aspects:
@@
    We will fix the bulk viscosity of the rotating solution to be the same as the nonrotating one, i.e., $\zeta=-\frac{1}{16\pi}$.
@@
Using the metric functions, the generic metric can be recast as
@@
Now, we should decompose this metric in the $(2+1+1)$ form:
@@
Let $u_{\mu}dx^{\mu}=F_{t}dt+\omega d\phi$, $n_{\mu}dx^{\mu}=F_{r}dr$. The structure of this metric can be put in the form of (44) such that one can directly start to calculate the factors relevant to the membrane description. The results for the relevant expressions are cumbersome; therefore we delegate these to the Appendix; here, we only give the pressure in the relevant limits. The pressure generically is of the form $P(r,\theta;\alpha,z,b,a,\epsilon_{3})$; and in the limit $z\rightarrow 0$ and $a\rightarrow 0$, it reduces to
which at the Schwarzschild horizon, becomes
So, this is the pressure obtained in the nonrotating limit of the JP metric: it depends on the deformation parameter $\epsilon_{3}$. On the other hand, in (39) we obtained an apparently different value for the surface gravity when we directly did the calculation in the static black hole metric. Therefore, ostensibly the no-rotation case and the limit of no rotation seem to yield different results. This is not acceptable; hence, we should identify these two relations, which necessarily fix one of the deformation parameters in terms of the other as
$\kappa_{\text{nonrotating}}\overset{!}{=}\kappa_{\text{doubly modified}}\implies\frac{1}{4m\left(\frac{\epsilon_{3}}{8}+1\right)}=\frac{1}{4m}\sqrt{\frac{1}{\frac{\alpha}{8}+1}},$
This identification also allows us to fix the value of the bulk viscosity to be $\zeta=-\frac{1}{16\pi}$, in the static limit of the JP black hole, while the shear viscosity is $\eta=\frac{1}{16\pi}$ and uniquely solve the transport coefficients for this case [22]. There are three things to note here: First, the two additional hairs $\alpha$ and $\epsilon_{3}$ are not independent; Second, the effective membrane description is valid only for the region $\epsilon_{3}>-8$. It is interesting to note that the critical point $\epsilon_{3}=-8$ was found before in a completely different context [9]. Finally the ratio $\frac{\zeta}{\eta}=-1\leq\frac{2}{3}$ is consistent with the bound computed from the AdS/CFT correspondence [27].
In Fig. 1, pressure versus the radial coordinate is plotted for all $r$ for the JP black hole. There is a discontinuity at the location of the ergosphere, and the pressure vanishes at the inner and outer horizons while diverging at the central singularity. It also asymptotically vanishes as $r\rightarrow\infty$. To compare we have also plotted as in Fig. 2 the pressure of the fluid for the Kerr black hole (see [22] where this plot was first given). The JP black hole and the Kerr black hole have similar $P$ vs $r$ graphs: The main difference is the sign of the pressure at the central singularity. In Fig. 3, we plotted the $P$ vs $r$ graph for the doubly modified Schwarzschild black hole. It is similar to the ordinary Schwarzschild case.
Figure 1: First a caution about this and the other figures; strictly speaking the membrane does not extend inside of the horizon. But it is remarkable to see that analytically continuing the pressure to the inside of the horizon, the fluid manages to capture the important surfaces of the black hole. For this reason, we plotted the pressure versus the radial coordinate over the whole range up to the central singularity. This figure represents the scaled pressure as a function of the radial coordinate $r$ for the choices $\alpha=1,m=1,b=0,a=0.75,\epsilon_{3}\to 6\sqrt{2}-8$. For these choices the Kerr horizon radii read as $r_{\text{H}_{outer}}=1.66$ while $r_{\text{H}_{inner}}=0.33$. One can see that at the ergosphere radius at the equator, $r_{\text{ergosphere}}=2$, there is a discontinuity in the pressure. One observes that the pressure diverges positively at the central singularity, discontinuous at the ergosphere radius at the equatorial plane, and asymptotically zero at infinity while vanishing at both the inner and outer event horizons.
Figure 3: This figure represents pressure values of the doubly modified Schwarzschild black hole scaled with $8\pi$ when we choose $m=1,a=0,\alpha=1$, the Schwarzschild radius at $r_{\text{H}}=2$. As we expected, the function is monotonically decreasing as $r$ gets larger reaching zero at asymptotic infinity, while positively diverging at the central singularity.
Now, let us check the transport coefficient that carries the energy flux into the null surface, i.e., the null expansion $\Theta$. When the metric coefficients are plugged in one has:
where $I(r,\theta)=2\left(a^{2}\cos^{2}\theta+r^{2}\right)^{2}$ and we have also included the $z$-deformation dependence. On the equatorial plane, in the nonrotating limit one gets
where we now can insert (49) to the null expansion while setting the deformation parameter $z=0$:
$\Theta_{|_{\alpha=0}}=2\left(1-\frac{2m}{r}\right),$ (53)
Equation (53) is the exact behavior found for the Schwarzschild black hole [22].
IV Conclusions and Discussions
Using the Parikh-Wilczek [15] action formulation, we developed a membrane description of the Johannsen-Psaltis black hole, which is a phenomenologically viable deformation of the Kerr metric that is amenable to test the Kerr hypothesis along with the no-hair theorem. As alluded to in the Introduction, a membrane is an effective description of a black hole as it appears to an outside observer. As such, it replaces the teleological concept of the black hole event horizon (a null hypersurface), which is not possible to probe by transient observers like us with a timelike fluid. The definition of the membrane is such that one might wonder if one is bartering the teleological event horizon with a tautological membrane that impersonates the black hole. This is not the case, but if it were the case, it would still not be so terrible because the membrane is local in time and has proved to provide an intuitive understanding of complicated phenomena such as the relativistic jet production by rotating black holes and their accretion disk. In this work, we have not studied the jet-production mechanism through the membrane paradigm, but the fact that one can construct a membrane for the JP metric for some deformation parameters is useful. Moreover, seemingly independent deformation parameters must be related for the membrane paradigm to work. We have ruled out some deformation parameters of the JP metric via the membrane paradigm. We have also shown that the ratio of the bulk viscosity to shear viscosity remains within the bounds predicted by the AdS/CFT dictionary.
As we have seen here the membrane can accommodate extra hairs yet constrain the values of these hairs. Our main task was to understand if a complicated metric such as the JP metric can fit into the membrane description, and we have seen that this is possible. If we analytically continue the pressure as a function of the radial coordinate $r$, we observe several interesting behaviors: At the location of the ergosphere on the equator, the pressure diverges just like the pressure of a Van der Waals gas diverges at the volume of the molecule. Therefore, the fluid is aware of the size of the ergoregion of the black hole. This might have interesting applications in astrophysical black holes. For example, in [28], it was shown that the Blandford-Znajek [29] that explains the relativistic jet production of rotating black holes with accretion discs is related to the ergosphere of the black hole and not to the event horizon. If the jet production is to be explained by the membrane paradigm it is clear that our construction above lends support to the computations of [28].
V Acknowledgments
We would like to thank Professor Dr. Sakir Erkoc for their support of this work.
VI Appendix: Details of Calculations for the Construction of the Membrane
The components of the extrinsic curvature of the metric (46) are found to be
@@ VI Appendix: Details of Calculations for the Construction of the Membrane
$K_{\theta\theta}$ $=\frac{\partial_{r}S}{2\sqrt{\frac{S}{\Delta Z}}},\quad K_{\phi\phi}=\frac{1}{E(r,\theta)}\left(\sin^{2}\theta S\left(\left(a^{2}+r^{2}\right)^{2}A^{2}-a^{2}\sin^{2}\theta\Delta\right)\right),\quad K_{rr}=0,$ (57)
where we introduced the following functions:
@@
Following the construction of a generic Kerr-like membrane paradigm algorithm, one can find the other transport coefficients. For instance, the nonzero components of the shear tensor turn out to be
@@
and the nonzero component of the momentum becomes
@@
When the metric coefficients are plugged to (64) it becomes